lilcutie2:

There are 22 adults and 19 kids in line for ice cream. The line is just too long and slow, so 8 people leave. How many people are in line now?

lilcutie2:

so 22 plus 19 =41 and then 41-8=33

TETSXPREMEE:

thats alot of adults imagine

nettym:

the adults are more desperate for ice cream then the kids lmao

hoezluvvjayy:

wait i read wrong lol

jesternmais:

19+22=41 41-8=33

jesternmais:

well it did n0t be specific with how many adults where taken out or kids so you don't need to watch that

jesternmais:

so your answer is 33